On a spectral notion of integrability

oliver knill harvard university

We discuss a spectral notion of integrability for dynamical systems. The definition applies to a large class of systems including systems with higher-dimensional time. We will look at the relation with other notions of integrability on classes of finite or infinite dimensional Hamiltonian systems.